On mr. Varum's most recent test, students were awarded 6 points for a correct response amd lost 3 points for incorrect responses. There were a total of 40 questions on the test.
.what would your score if you answered 34 questions correctly?
.what would your score be if you answered 19 questions incorrectly?

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

x the number of correct answers 40-x incorrect
6x - 3(40-x) = Score
x = 34
6*34 -3(40-34) = 204 - 3(6) = 186

Answer by stanbon(75887)

On mr. Varum's most recent test, students were awarded 6 points for a correct response amd lost 3 points for incorrect responses. There were a total of 40 questions on the test.
.what would your score if you answered 34 questions correctly?
Ans: 6*34 - 3*6 = 186 points
-----------------------------
.what would your score be if you answered 19 questions incorrectly?
Ans: 6(21) - 3*19 = 69 points
--------------------
Cheers,
Stan H.
------------------